# CARTOS: A Charging-Aware Real-Time Operating System for Intermittent Batteryless Devices

### v2

@@ III-B1 Mixed-Preemption Model
III-B2 Support for Periodic Tasks and Chains
@@
Fig. 5 depicts the example of a periodic task (task1Runner) in CARTOS. setTaskPeriod() is called at the beginning of the task’s code, declaring that it is a periodic task. The task has a while-loop where each iteration represents a job for each period. When the task reaches the end of execution in the current period, it calls waitForNextPeriod(). Then, the scheduler removes it from the ready queue and sets the system timer to insert the task back to the ready queue at the start of the next period.
@@ III-C Adaptive Real-Time Scheduler
If the task $\tau_{h}$ is non-atomic (computational), the scheduler executes $\tau_{h}$ for one tick of time (scheduling time quantum; 1 ms in our implementation) and repeats the above procedure to find the highest-priority ready task again for preemptiveness. Note that this is the case for tick-based RTOS. In the case of a tickless system, $\tau_{h}$ can continue to run until it voluntarily relinquishes the CPU or any higher-priority task than $\tau_{h}$ becomes ready.
If $\tau_{h}$ is atomic (peripheral), the scheduler checks the required energy for $\tau_{h}$’s job execution based on the profiled data, and compares it with the current system energy level, measured through the voltage level of the capacitor (more details on this will be given in the next subsection). If the system’s available energy is higher than the amount required to ensure the uninterrupted execution of the given atomic task, the scheduler dispatches this task and the task will continue until its job is completed, i.e., non-preemptive execution. Otherwise, it proactively enforces a power cycle to harvest more energy. [^3] Precisely speaking, the device transitions to the standby phase using the energy controller unit discussed in Sec. III-A. In doing so, it first estimates the required harvesting time, $\Delta t$, considering the predicted charging rate and the current voltage level, [^4] It is worth noting that the harvesting time $\Delta t$ is not necessarily equal to the maximum charging demand of the task ($Q_{i}^{j}$, explained in Sec III-E) because the device’s current voltage can be higher than the low-voltage threshold. and then computes the next device wake-up time from the standby phase by taking into account both the estimated harvesting time and the future release times of other higher-priority periodic tasks. This is done to ensure that higher-priority tasks are not missed during charging. For example, if a high-priority task $\tau_{1}$ needs to be released at time $t_{1}$ while the device is being charged for a low-priority task $\tau_{2}$, the charging process should be stopped so that the high-priority task $\tau_{1}$ is scheduled. Therefore, the next wake-up time from the standby phase is obtained by taking the minimum between the required harvesting time and the earliest future release time of higher-priority tasks than $\tau_{h}$.
Subsequently, the device is powered down (after necessary JIT checkpoint operations), and the atomic task $\tau_{h}$ as well as other tasks are scheduled when the device is powered on. It should be noted that such a proactive power-cycling sequence happens before reaching the low-voltage threshold. When the device reaches a low-voltage threshold, the device power cycling is managed by the JIT checkpointing service presented next.
@@ III-D JIT Checkpointing Service
In CARTOS, JIT checkpointing is implemented as a system service task. This service monitors if the capacitor voltage level reaches the low-voltage threshold, which can be implemented as an external interrupt or periodic polling of the voltage monitoring circuits. The JIT service task is launched as part of the boot-up sequence and executes with the highest priority in the system to execute as soon as possible when needed.
The low-voltage threshold can be triggered at any time during non-atomic preemptible task execution (recall that our scheduler executes atomic tasks only when the device has sufficient energy to complete their jobs). Then, the JIT service task enforces a new power cycle in a similar procedure as stated before. Specifically, the next wake-up time is calculated as the minimum of the earliest future release time of other higher-priority tasks and the estimated harvesting time $\Delta t$ required to complete the ‘remaining portion’ of the currently running task. Next, it performs JIT checkpointing operations. It iterates over all non-atomic tasks in the system, and for each non-atomic task, it stores the task’s context (task control block), stack, and SRAM data areas to NVM so that those can be restored to SRAM upon the next power-on. The JIT service also stores the current system time and timer events in NVM in order to ensure the correct execution of timer-based operations across power cycles. Once the device powers up, the boot-up procedure restores non-atomic tasks and updates the system time and timer events. The time update can be done in two ways. If there is an external RTC, e.g., as part of the energy controller, it can simply update the system time based on it. Otherwise, it can estimate the current time by adding the latest checkpointed system time and the power-down time that was computed before power cycling (Sec. III-C and Fig. 6). We use the latter approach in our implementation.
@@ III-E Charging Management
While JIT checkpointing facilitates the proper execution of non-atomic tasks, charging management is vital to ensure the execution of atomic tasks. To address this, we assign a voltage threshold $V^{j}_{i}$ to each atomic task $\tau^{j}_{i}$. This threshold represents the minimum voltage that the system requires before starting the execution of task $\tau^{j}_{i}$’s job, thereby ensuring that the task can be completed before reaching the low-voltage threshold used for JIT checkpointing. The calculation for $V^{j}_{i}$ is as follows:
@@
Eq. (1) is derived from the well-known equation [41] on the energy stored in capacitor when it receives a fixed power of $W$ from time $t_{1}$ to $t_{2}$, given by
@@
In Eq. (1), the values of $C^{j}_{i}$ and $Q^{j}_{i}$ are specific to each task, while $V_{min}$ and $\mathbb{C}$ are the system parameters given by the user at design time. These parameters should be designed so that for all atomic tasks, $V_{i}^{j}\leq V_{max}$ where $V_{max}$ is the maximum voltage that can be reached by the given energy harvester and the capacitor’s nominal voltage. The charging rate of the system, $W_{S}$, is updated at runtime using an energy predictor as the environment changes. Additionally, after each job execution, the value of the worst-case execution time $C^{j}_{i}$ is profiled and updated in the task’s task control block (TCB). If there is a change in $C^{j}_{i}$ or $W_{S}$, the value of $Q^{j}_{i}$ for that task is also updated at runtime. The calculation for updating $Q^{j}_{i}$ is as follows:
@@
$\Delta t=\frac{\mathbb{C}\left({V^{j}_{i}}^{2}-{V_{current}}^{2}\right)}{2\cdot W_{S}}$ (4)
On the other hand, when scheduling non-atomic tasks, the scheduler does not have to consider the current voltage, as depicted in Fig. 6. Hence, as long as the current voltage is above $V_{min}$, a non-atomic task can begin execution. However, if the voltage drops to or below $V_{min}$ during the execution of the non-atomic task, the JIT checkpointing task is activated and computes the required harvesting time $\Delta t$ using Eq. (4). Specifically, $V_{i}^{j}$ is computed for the remaining portion of the task and capped by $V_{max}$ so that a long-running non-atomic task can execute over multiple power cycles. It is worth noting that our scheduler design improves the average-case runtime performance of non-atomic tasks by preventing unnecessary blocking and enables reclaiming unused charges from other tasks. However, in the worst case, the maximum total delay experienced by the task is still up to $Q_{i}^{j}$ by definition.
The energy harvesting rate, $W_{S}$, can change over time depending on environmental conditions. As the required harvesting time and the next device wake-up time rely on this, it is crucial to obtain a precise value of $W_{S}$ for efficient operations because the over-estimation makes the device turn on unnecessarily late and the under-estimation lets the device turn on and off multiple times until it reaches the required charge level. To address this issue, CARTOS employs a machine learning-based energy prediction model proposed in [24]. This model is a simple yet viable approach for MCU-level devices, providing a reliable prediction of solar harvesting rates by leveraging the temporal dependencies of time-series voltage data. In our implementation, we integrated this model into CARTOS and used a 30-minute interval for our prediction window.
@@ IV-A System Model
We consider a system with $n$ processing chains. A chain $i$ is modeled as $\Gamma_{i}:(C_{i},T_{i},D_{i},Q_{i})$, where $C_{i}$, $T_{i}$, $D_{i}$, and $Q_{i}$ are the cumulative worst-case execution time, period, deadline, and total charging demand of one instance (job) of the chain, i.e., $C_{i}=\sum_{j=1}^{m_{i}}C^{j}_{i}$ and $Q_{i}=\sum_{j=1}^{m_{i}}\max(Q^{j}_{i},0)$. The max term is used to handle a negative value for $Q^{j}_{i}$, which can occur when the discharging rate of a task is lower than the energy-harvesting rate of the system. All chains have constrained deadlines, i.e., $\forall i:D_{i}\leq T_{i}$. Each chain $\Gamma_{i}$ is comprised of $m_{i}$ tasks. Each task is either atomic (non-preemtable) or non-atomic (preemptible), following the mixed-preemption model of CARTOS. A task $j$ of a chain $\Gamma_{i}$ is denoted as $\tau^{j}_{i}:(C^{j}_{i},T^{j}_{i},D^{j}_{i},Q^{j}_{i})$, where $\forall j\leq m_{i}~{}|~{}T^{j}_{i}=T_{i}$, i.e., all the tasks within the same chain share the same period as their chain and the execution time of the chain is the sum of the execution of all tasks in that chain. Tasks in a processing chain are ordered in sequence. Hence, for a processing chain to complete, all of its tasks should complete their execution in that order, i.e. $\tau^{1}_{i}$ should be the first task to be executed and then $\tau^{2}_{i}$ and so on.
@@
In real-world scenarios, the energy harvesting rate of the system and the charging demand of tasks can vary with the availability of energy resources. However, especially for the solar energy we consider in this work, the changes are relatively slow (e.g., over several tens minutes and hours) and do not typically happen within a chain period or the hyperperiod of chains. Therefore, a fixed energy harvesting rate is considered for analysis purposes while our runtime system can still adapt to such changes.
@@ IV-B Schedulability Analysis
Recall our mixed-preemption scheduling model. Atomic tasks are non-preemptively scheduled, so once started, they cannot be preempted by any other tasks. On the other hand, non-atomic tasks are preemptible by higher-priority tasks during their execution. To analyze the worst-case response time of a chain comprised of such tasks, we extend two existing methods: the task-level mixed preemption analysis without energy constraint [42], and the non-preemptible IPD task scheduling analysis [13], both of which are exact analysis. Our extension takes into account the charging requirements of tasks and the execution order within chains, as follows.
@@
[…]
@@
where $A_{l}^{j}$ indicates the atomicity of a task $\tau_{l}^{j}$, i.e., $A_{l}^{j}=true$ if $\tau_{l}^{j}$ is atomic (non-preemptible) and $A_{l}^{j}=false$ if $\tau_{l}^{j}$ is non-atomic (preemptible). The iteration of Eq. (5) starts with $L_{i}^{\left(0\right)}=B_{i}+C_{i}$ and stops when $L_{i}^{\left(s\right)}=L_{i}^{\left(s-1\right)}$ (converged) or $L_{i}^{\left(s\right)}\geq T_{H}$ where $T_{H}$ is the hyperperiod of all chains in the system (failed). Here, the charging demand $Q_{h}$ is added to the execution time $C_{h}$ because CARTOS can effectively delay execution by up to this amount when the charge is insufficient. Recall our scheduler and charging management in Sec. III. The wake-up time is given as the minimum between the harvesting time $\Delta t$ and the earliest release time, and $Q_{i}^{j}$ is greater than or equal to $\Delta t$ by definition (Sec. III-E). Hence, using $Q_{h}$ in the analysis provides a safe upper bound for all charging-related delays of individual tasks of $\Gamma_{h}$ in the worst case.
@@
[…]
To find the start time of task $\tau_{i,k}^{m_{i}}$, $S_{i,k}^{m_{i}}$, we should consider the following factors: (i) blocking caused by lower priority chains, (ii) preemption caused by higher priority chains, (iii) execution time of previous jobs of the chain $\Gamma_{i}$, (iv) executions of all other tasks in the chain $\Gamma_{i}$ that should be serviced before the task $\tau_{i,k}^{m_{i}}$, and (v) all the overheads caused by charging. Note that the job of the chain $\Gamma_{i}$ can be blocked only once during its entire job execution due to the priority-based scheduling of CARTOS. Therefore, for the task $\tau_{i,k}^{m_{i}}$, the start time $S_{i,k}^{m_{i}}$ can be computed by the following recurrence:

### v3

@@ III-B2 Memory Model
III-B3 Support for Periodic Tasks and Chains
@@
Fig. 5 depicts the example of a periodic task in CARTOS. setTaskPeriod() is called at the beginning, declaring that it is a periodic task. The task has a while-loop where each iteration represents a job for each period. When the task reaches the end of execution in the current period, it calls waitForNextPeriod(). Then, the scheduler removes it from the ready queue and sets the system timer to insert the task back to the ready queue at the start of the next period.
@@ III-C Adaptive Real-Time Scheduler
If $\tau_{h}$ is non-atomic (computational), the scheduler executes $\tau_{h}$ for one tick of time (scheduling time quantum; 1 ms in our implementation) and repeats the above procedure to find the highest-priority ready task again for preemptiveness. Note that this is the case for tick-based RTOS. In the case of a tickless system, $\tau_{h}$ can continue to run until it voluntarily relinquishes the CPU or is preempted by a higher-priority task.
If $\tau_{h}$ is atomic (peripheral), the scheduler checks the required energy for $\tau_{h}$’s job execution based on the profiled data, and compares it with the current system energy level, measured through the voltage level of the capacitor (details in Sec. III-E). If the system’s available energy is higher than the amount required to ensure the uninterrupted execution of the given atomic task, the scheduler dispatches this task and the task will continue until its job is completed, i.e., non-preemptive execution. Otherwise, it proactively enforces a power cycle to harvest more energy. [^4] Precisely speaking, the device transitions to the standby phase using the energy controller unit discussed in Sec. III-A. In doing so, it first estimates the required harvesting time, $\Delta t$, considering the predicted charging rate and the current voltage level, [^5] It is worth noting that the harvesting time $\Delta t$ can be smaller than the maximum charging demand of the task ($Q_{i}^{j}$, explained in Sec III-E) because the device’s current voltage can be higher than the low-voltage threshold. and then computes the next device wake-up time from the standby phase by taking into account both the estimated harvesting time and the future release times of other higher-priority periodic tasks. This is done to ensure that higher-priority tasks are not missed during charging. For example, if a high-priority task $\tau_{1}$ needs to be released at time $t_{1}$ while the device is being charged for a low-priority task $\tau_{2}$, the charging process should be stopped so that the high-priority task $\tau_{1}$ is scheduled. Therefore, the next wake-up time from the standby phase is obtained by taking the minimum between the required harvesting time and the earliest future release time of higher-priority tasks than $\tau_{h}$.
Subsequently, the device is powered down (after necessary JIT checkpoint operations for preempted non-atomic tasks). The atomic task $\tau_{h}$ as well as other tasks are scheduled when the device is powered on. It should be noted that such a proactive power-cycling sequence happens before reaching the low-voltage threshold. When the device reaches a low-voltage threshold, the device power cycling is managed by the JIT checkpointing service presented next.
@@ III-D JIT Checkpointing Service
In CARTOS, JIT checkpointing is implemented as a system service task. This service monitors if the capacitor voltage level reaches the low-voltage threshold, which can be implemented as an external interrupt or periodic polling of the voltage monitoring circuits. The JIT service task is launched as part of the boot-up sequence and executes with the highest priority to execute as soon as possible when needed.
JIT checkpointing occurs only once per power cycling (not every scheduling tick), which happens in two cases as depicted in Fig. 6: (i) proactive power cycling before atomic task execution (“Is energy enough $\rightarrow$ No” path in the figure), and (ii) reactive power cycling by the low-voltage threshold (“Low voltage threshold triggered” path). In both cases, the next device wake-up time needs to be computed. Since its calculation in case (i) has been discussed already, we now explain that in case (ii). Recall that the low-voltage threshold is triggered only during non-atomic task execution (since atomic tasks execute only with sufficient energy). Let $\tau_{i}$ denote this interrupted non-atomic task. To compute the next device wake-up time, the system needs $\Delta t$ to harvest energy for $\tau_{i}$’s remaining portion. However, higher-priority (HP) tasks may be released during charging. To ensure these HP tasks can preempt $\tau_{i}$ promptly, the system wakes up at the minimum of: (i) the earliest release time of HP tasks, and (ii) the harvesting time $\Delta t$ for $\tau_{i}$’s completion. Upon wake-up, the scheduler follows Fig. 6 to handle the highest-priority ready task at the moment.
Once the next device wake-up time is computed, JIT checkpointing is performed. It iterates over all non-atomic tasks in the system, and for each non-atomic task, it stores the task’s context (task control block), stack, and SRAM data areas (if exist) to NVM so that those can be restored to SRAM upon the next power-on. The JIT service also stores the current system time and timer events in NVM in order to ensure the correct execution of timer-based operations across power cycles.
When the device powers up, the boot-up procedure restores non-atomic tasks and updates the system time and timer events. The time update can be done in two ways. If there is an external RTC, e.g., as part of the energy controller, it can simply update the system time based on it. Otherwise, it can estimate the current time by adding the latest checkpointed system time and the programmed power-down duration (which equals the computed device wake-up time). We use the latter approach in our implementation.
If the low-voltage threshold turns out to be insufficient to complete JIT checkpointing before power loss, the system will detect an incomplete or invalid checkpoint upon next reboot. In this case, affected non-atomic tasks simply restart from their beginning, as if they were executing for the first time, and subsequent scheduling operations are performed normally.
@@ III-E Charging Management
Charging management is vital to ensure the correct execution of atomic tasks. To address this, we assign a voltage threshold $V^{j}_{i}$ to each atomic task $\tau^{j}_{i}$. This threshold represents the minimum voltage that the system requires before starting the execution of task $\tau^{j}_{i}$’s job, thereby ensuring that it can complete before reaching the low-voltage threshold used for JIT checkpointing. The calculation for $V^{j}_{i}$ is as follows:
@@
Eq. (1) is derived from the well-known equation [halliday2013] on the energy stored in capacitor when it receives a fixed power of $W$ from time $t_{1}$ to $t_{2}$, given by
@@
In Eq. (1), the values of $C^{j}_{i}$ and $Q^{j}_{i}$ are specific to each task, while $V_{min}$ and $\mathbb{C}$ are the system parameters given by the user at design time. These parameters should be designed so that for all atomic tasks, $V_{i}^{j}\leq V_{max}$ where $V_{max}$ is the maximum voltage that can be reached by the given energy harvester and the capacitor’s nominal voltage. The charging rate of the system, $W_{S}$, represents a conservative minimum among observations taken at short intervals (e.g., every second, as in [Celebi2020, Zygarde, MohsenIOTJ2021]). The worst-case execution time $C^{j}_{i}$ (obtained a priori) is stored in the TCB and can be updated at runtime to deal with potential underestimation. If there is a change in $C^{j}_{i}$ or $W_{S}$, the value of $Q^{j}_{i}$ for that task is also updated at runtime. The calculation for updating $Q^{j}_{i}$ is as follows:
@@
$\Delta t=\frac{\mathbb{C}\left((\min{({V^{j}_{i}},V_{max})})^{2}-{V_{current}}^{2}\right)}{2\cdot W_{S}}$ (4)
Here, $V_{i}^{j}\leq V_{max}$ holds for atomic tasks, as discussed above.
On the other hand, when scheduling non-atomic tasks, the scheduler does not have to consider the current voltage, as depicted in Fig. 6. Hence, as long as the current voltage is above $V_{min}$, a non-atomic task can begin execution. However, if the voltage drops to or below $V_{min}$ during the execution of the non-atomic task, the JIT checkpointing task is activated and computes the required harvesting time $\Delta t$ using Eq. (4). Specifically, $V_{i}^{j}$ is computed for the remaining portion of the task and capped by $V_{max}$ (i.e., $\min{({V^{j}_{i}},V_{max})}$ in Eq. (4)) so that a long-running non-atomic task can execute over multiple power cycles. It is worth noting that our scheduler design improves the average-case runtime performance of non-atomic tasks by reclaiming unused charges from other tasks. However, in the worst case, the maximum total delay experienced by the task is still up to $Q_{i}^{j}$ by definition.
The energy harvesting rate, $W_{S}$, can change over time depending on environmental conditions. As the required harvesting time and the next device wake-up time rely on this, it is crucial to obtain a precise value of $W_{S}$ for efficient operations because the over-estimation makes the device turn on unnecessarily late and the under-estimation lets the device turn on and off multiple times until it reaches the required charge level. To address this issue, CARTOS employs a machine learning-based energy prediction model proposed in [MohenRTCSA2022]. This model is a simple yet viable approach for MCU-level devices, providing a reliable prediction of solar harvesting rates by leveraging the temporal dependencies of time-series voltage data. In our implementation, we integrated this model into CARTOS and used a 30-minute interval for our prediction window.
Using a single conservative harvesting rate $W_{S}$ within each prediction interval may lead to pessimism in highly dynamic environments. To reduce this pessimism, one may consider more sophisticated approaches, e.g., [Zygarde] captures the probability of $N$ observations of a specific energy harvesting rate, but this is beyond the scope of our paper.
@@ IV-A System Model
We consider a system with $n$ processing chains. A chain $i$ is modeled as $\Gamma_{i}:(C_{i},T_{i},D_{i},Q_{i})$, where $C_{i}$, $T_{i}$, $D_{i}$, and $Q_{i}$ are the cumulative worst-case execution time, period, deadline, and total charging demand of one instance (job) of the chain, i.e., $C_{i}=\sum_{j=1}^{m_{i}}C^{j}_{i}$ and $Q_{i}=\sum_{j=1}^{m_{i}}\max(Q^{j}_{i},0)$. The max term is used to handle a negative value for $Q^{j}_{i}$, which can occur when the discharging rate of a task is lower than the energy-harvesting rate of the system. All chains have constrained deadlines, i.e., $\forall i:D_{i}\leq T_{i}$. Each chain $\Gamma_{i}$ is comprised of $m_{i}$ tasks. Each task is either atomic (non-preemtable) or non-atomic (preemptible), following the mixed-preemption model of CARTOS. A task $j$ of a chain $\Gamma_{i}$ is denoted as $\tau^{j}_{i}:(C^{j}_{i},T^{j}_{i},D^{j}_{i},Q^{j}_{i})$, where $\forall j\leq m_{i}~|~T^{j}_{i}=T_{i}$, i.e., all the tasks within the same chain share the same period as their chain and the execution time of the chain is the sum of the execution of all tasks in that chain. Tasks in a processing chain are ordered in sequence. Hence, for a processing chain to complete, all of its tasks should complete their execution in that order, i.e. $\tau^{1}_{i}$ should be the first task to be executed and then $\tau^{2}_{i}$ and so on. It is important to note that all atomic tasks should satisfy: $V_{i}^{j}\leq V_{max}$, as explained in Sec. III-E. For non-atomic tasks, we assume that the low-voltage threshold is sufficient to complete JIT checkpointing before power loss. The timing guarantees from our analysis do not hold when these assumptions are violated. Nonetheless, the analysis is useful before deployment as it provides insight into the system’s behavior under the assumed conditions.
@@
In real-world scenarios, the energy harvesting rate of the system and the charging demand of tasks can vary with the availability of energy resources. However, especially for the solar energy we consider in this work, the changes are relatively slow (e.g., over several tens of minutes and hours) and do not typically happen within a chain period or the hyperperiod of chains. Therefore, a fixed energy harvesting rate is considered for analysis purposes while our runtime system can still adapt to such changes.
@@ IV-B Schedulability Analysis
Recall our mixed-preemption scheduling model. Atomic tasks are non-preemptively scheduled, so once started, they cannot be preempted by any other tasks. On the other hand, non-atomic tasks are preemptible by higher-priority tasks during their execution. To analyze the worst-case response time of a chain comprised of such tasks, we extend two existing methods: the task-level mixed preemption analysis without energy constraint [regehr2002scheduling], and the non-preemptible IPD task scheduling analysis [MohsenIOTJ2021], both of which are exact analysis. Our extension takes into account the charging requirements of tasks and the execution order within chains, as follows.
@@
where $B_{i}$ is the blocking time due to a non-preemptible task from a lower-priority chain, $Q_{h}$ is the cumulative charging demand of a high-priority chain $\Gamma_{h}$, and $\left\lceil\right\rceil$ is the ceiling function. $B_{i}$ can be obtained by
@@
where $A_{l}^{j}$ indicates the atomicity of a task $\tau_{l}^{j}$, i.e., $A_{l}^{j}=true$ if $\tau_{l}^{j}$ is atomic (non-preemptible) and $A_{l}^{j}=false$ if $\tau_{l}^{j}$ is non-atomic (preemptible). The iteration of Eq. (5) starts with $L_{i}^{\left(0\right)}=B_{i}+C_{i}$ and stops when $L_{i}^{\left(s\right)}=L_{i}^{\left(s-1\right)}$ (converged) or $L_{i}^{\left(s\right)}\geq T_{H}$ where $T_{H}$ is the hyperperiod of all chains in the system (failed). Here, the charging demand $Q_{h}$ is added to the execution time $C_{h}$ because CARTOS can effectively delay execution by up to this amount when the charge is insufficient. Recall our scheduler and charging management in Sec. II-D. The wake-up time is given as the minimum between the harvesting time $\Delta t$ and the earliest release time, and $Q_{i}^{j}$ is greater than or equal to $\Delta t$ by definition (Sec. III-E). Hence, using $Q_{h}$ in the analysis provides a safe upper bound for all charging-related delays of individual tasks of $\Gamma_{h}$ in the worst case.
@@
where $\left\lfloor\right\rfloor$ is the floor function. The iteration in Eq. (10) can start with ${F_{i,k}}^{(0)}=S_{i,k}^{m_{i}}+C^{m_{i}}_{i}$.
To find the start time of task $\tau_{i,k}^{m_{i}}$, $S_{i,k}^{m_{i}}$, we should consider the following factors: (i) blocking caused by lower priority chains, (ii) preemption caused by higher priority chains, (iii) execution time of previous jobs of the chain $\Gamma_{i}$, (iv) executions of all other tasks in the chain $\Gamma_{i}$ that should be serviced before the task $\tau_{i,k}^{m_{i}}$, and (v) all the extra time caused by charging. Note that the job of the chain $\Gamma_{i}$ can be blocked only once during its entire job execution due to the priority-based scheduling of CARTOS. Therefore, for the task $\tau_{i,k}^{m_{i}}$, the start time $S_{i,k}^{m_{i}}$ can be computed by the following recurrence: